a fruit company guarentees that 90% of the pineapples it ships will be ripe within four days. find each probability for a case containing 12 pineapples.
1) all 12 are ripe within 4 days.

2) at least 10 are ripe within 4 days.

3) no more than 9 are ripe within 4 days.

Answer by stanbon(75887)

a fruit company guarentees that 90% of the pineapples it ships will be ripe within four days. find each probability for a case containing 12 pineapples.
---
Binomial Problem with n=12 ; p = 0.90
-------
1) all 12 are ripe within 4 days.
P(x=12) = 12C12*(0.9)^12*(0.1)^0 = 0.9^12 = 0.2824
---------------------------------------------
2) at least 10 are ripe within 4 days.
P(10<= x <=12) = 1 - binomcdf(12,0.9,9) = 0.8891
---------------------------------------------
3) no more than 9 are ripe within 4 days.
P(0<= x <=9) = binomcdf(12.0.9,9) = 0.1109
=============================================
Cheers,
Stan H.